# Auditing Political Exposure Bias: Algorithmic Amplification on Twitter/$\mathbb{X}$ Approaching the 2024 U.S. Presidential Election

### v2

@@
Abstract
Approximately 50% of tweets in $\mathbb{X}$’s user timelines are personalized recommendations from accounts they do not follow. This raises a critical question: what political content are users exposed to beyond their established networks, and how might this influence democratic discourse online? Due to the black-box nature and constant evolution of social media algorithms, much remains unknown about this aspect of users’ content exposure, particularly as it pertains to potential biases in algorithmic curation. Prior research has shown that certain political groups and media sources are amplified within users’ in-network tweets. However, the extent to which this amplification affects out-of-network recommendations remains unclear. As the 2024 U.S. Election approaches, addressing this question is essential for understanding the influence of algorithms on online political content consumption and its potential impact on users’ perspectives.
@@ Abstract
In this paper, we conduct a three-week audit of $\mathbb{X}$’s algorithmic content recommendations using a set of 120 sock-puppet monitoring accounts that capture tweets in their personalized “For You” timelines. Our objective is to quantify out-of-network content exposure for right- and left-leaning user profiles and to assess any potential biases in political exposure. Our findings indicate that $\mathbb{X}$’s algorithm skews exposure toward a few high-popularity accounts across all users, with right-leaning users experiencing the highest level of exposure inequality. Both left- and right-leaning users encounter amplified exposure to accounts aligned with their own political views and reduced exposure to opposing viewpoints. Additionally, we observe a right-leaning bias in exposure for new accounts within their default timelines. We hope this empirical study promotes discussion on the transparency and accountability of social media algorithms, contributing to critical issues on safeguarding election integrity and fostering a more informed digital public sphere.
Introduction
@@
During the 2024 U.S. Presidential Election, social media platforms like $\mathbb{X}$ (formerly Twitter) serve as key venues for political information and discourse. Meanwhile, the information users encounter on the $\mathbb{X}$ platform is increasingly curated by algorithmic recommendation systems that personalize content in their “For You” timelines. As of this writing, $\mathbb{X}$’s “For You” timeline typically consists of 50% in-network tweets (i.e., from accounts a given user follows) and 50% out-of-network tweets (i.e., from accounts that user does not directly follow)—up from 37% in 2023 (Bouchaud et al., 2023). How does $\mathbb{X}$’s algorithm select relevant tweets from outside a user’s network? In 2023, Twitter partially open-sourced its recommendation algorithm, revealing that out-of-network recommendations are sourced through engagement and follow graphs, ranked by a neural network, and refined with heuristics and filters (Twitter, Inc., 2023). Despite this algorithm’s growing influence, little has been done to examine the specific composition and nature of these out-of-network tweets.
The impact of algorithmic content curation on political discourse in social media has been a major focus of research and public debate: Previous studies consistently show that $\mathbb{X}$’s algorithm amplifies political biases and prioritizes high-engagement content, including emotionally charged, toxic, and low-credibility information (Huszár et al., 2022; Bouchaud et al., 2023; Bandy and Diakopoulos, 2021; Chen et al., 2020; Bartley et al., 2021; Corsi, 2024). Researchers have used methods including randomized experiments, sock-puppet audits, crowdsourced audits, and observational data to study $\mathbb{X}$’s algorithmic effects. Some have found that Twitter’s algorithms tend to amplify content from right-leaning media sources and politicians more than their left-leaning counterparts (Huszár et al., 2022). Other studies report increased exposure to ideologically aligned friends (Bouchaud et al., 2023; Bandy and Diakopoulos, 2021) and low-credibility content (Corsi, 2024), with right-leaning users experiencing higher exposure to such content (Chen et al., 2020). For out-of-network tweets, qualitative analyses suggest the algorithm leans toward promoting centrist content for partisan users (Chen et al., 2020) and displays a more diverse political mix overall (Bouchaud et al., 2023).
@@
Despite these insights, current research on $\mathbb{X}$’s algorithmic auditing faces a critical challenge in analyzing out-of-network content: Many studies assess amplification by comparing personalized timelines with reverse-chronological timelines as baselines, where tweets appear in the order they were posted without algorithmic effects (Bartley et al., 2021, 2023; Huszár et al., 2022; Bouchaud et al., 2023). However, out-of-network tweets lack a reverse-chronological baseline, as users do not follow the authors of those tweets, making it challenging to quantitatively measure exposure bias. To address this limitation, we utilize a “sock-puppet audit,” a study design that deploys artificial user accounts with controlled features to systematically capture and analyze platform recommendations. This approach is particularly well-suited to studying out-of-network exposure patterns because it allows us to observe algorithmic behavior without the interference of real user behaviors or connections (Bartley et al., 2021; Chen et al., 2020; Bandy and Diakopoulos, 2021).
In this study, we deploy 120 sock-puppet accounts distributed across four groups—left-leaning, right-leaning, balanced, and neutral—enabling us to collect a robust dataset that currently encompasses over 5 million tweets. This large-scale audit offers a unique perspective on algorithmic content exposure, as it allows for comprehensive comparisons across various political profiles. Further details are provided in the Experimental Setup section.
@@
    We quantify algorithmic exposure to out-of-network content for users with varying political alignments during the 2024 U.S. election period through a sock-puppet audit of $\mathbb{X}$’s personalized timelines.
@@
We find that:
    The $\mathbb{X}$ platform skews exposure toward a few high-popularity accounts for all users, with right-leaning users experiencing the most inequality.
    Both left- and right-leaning users encounter amplified exposure to accounts aligned with their own political stance and reduced exposure to opposing viewpoints.
    Additionally, neutral accounts who do not follow anybody (akin to a newly-registered user account on the platform) show a default right-leaning bias in content exposure.
Experimental Setup
To analyze algorithmic biases in $\mathbb{X}$’s “For You” timeline, we developed a timeline scraper to systematically collect tweets recommended to different types of user profiles. We created 120 accounts divided into four groups: 30 neutral accounts (default setting, following no one), 30 left-leaning accounts, 30 right-leaning accounts, and 30 balanced accounts.
To categorize the political alignment of accounts to follow, we used the AllSides Media Bias Chart, [^1] AllSides Media Bias Chart https://www.allsides.com/media-bias/media-bias-chart which rates news sources on a spectrum from left to right based on their political bias. Each left-leaning and right-leaning account follows 10 media outlets, including seven outlets with a moderate (center-left or center-right) bias and three with a stronger (left or right) bias, as defined by the AllSides’ chart. This selection ensures that these accounts represent a realistic mix of moderately and strongly aligned sources, enhancing the accuracy of our analysis of political exposure. Additionally, left-leaning accounts follow key Democratic figures and entities (Kamala Harris, Tim Walz, House Democrats, and Senate Democrats), while right-leaning accounts follow their Republican counterparts (Donald Trump, JD Vance, House Republicans, and Senate Republicans). Balanced accounts, designed to reflect a centrist perspective, follow five center-left and center-right media outlets and both presidential candidates from each major party. All media follows are randomly selected from the respective groups in the media bias chart, ensuring consistency with each group’s intended alignment.
The timelines for each account are scraped four times daily, yielding approximately 500–700 tweets per session, or about 2,000–3,000 tweets per account per day, within the limits that $\mathbb{X}$’s terms of service impose on new, non-premium accounts. The choice of four daily scraping sessions was made to capture the variability in recommendations throughout the day, as the content recommended by $\mathbb{X}$’s algorithm can shift based on temporal factors like recent events or trending topics. This frequency provides a more comprehensive picture of the algorithmic exposure that users might experience. Data collection began on October 2, 2024, and continued through October 28, 2024. Regarding the proportion of out-of-network tweets, neutral accounts, which follow no other accounts, exclusively receive out-of-network content, while other account types have approximately 56–59% of their content coming from out-of-network sources. Further details on data collection are provided in the Methods section.
Results
Out-of-Network Exposure Inequality Among Users with Different Political Leanings
@@ Results
One significant aspect of algorithmic biases on social media is popularity bias (Nikolov et al., 2019). Algorithms often tend to amplify content from certain users over others, creating inequalities in exposure (Bartley et al., 2023). For instance, Twitter’s ranking algorithm employs a ~48M parameter neural network, which uses thousands of features to score each tweet based on engagement probabilities, prioritizing content with higher likelihoods of interaction in users’ feeds (Twitter, Inc., 2023).
@@
Table 1 provides an overview of the statistics for the collected tweet dataset across different account types. It shows the average proportion of out-of-network tweets that each account type encounters, with neutral accounts seeing exclusively out-of-network content, while the other accounts have around 56-59% of their timelines composed of out-of-network tweets. Additionally, it details the average proportions of retweets, quoted tweets, and promoted tweets observed by each account type.
@@
Table 1. Statistics of the collected tweet dataset (mean values with standard deviations)
@@
Out-of-network tweet 100% 56.51% (6.52) 57.41% (7.25) 59.00% (5.56)
Retweet 0.04% (0.11) 2.81% (1.02) 2.33% (1.30) 2.57% (1.86)
Quoted tweet 0.94% (1.56) 7.93% (2.12) 12.75% (3.33) 12.40% (2.59)
Promoted tweet 0.95% (1.44) 6.42% (1.09) 5.67% (1.14) 7.54% (0.62)
Previous research has shown that popularity biases can lead to a skew in the visibility of tweets when comparing personalized feeds with reverse-chronological ones, and that users are disproportionately exposed to friends’ tweets (Bartley et al., 2023, 2021). Yet, it remains unclear whether exposure inequalities extend beyond friends to include a broader set of recommended users. Specifically, we ask:
Are personalized recommendations in $\mathbb{X}$ distributed evenly among users, or is exposure dominated by a few accounts? Furthermore, do these inequalities differ among users with different political leanings?
@@
To address these research questions, we use the Gini coefficient, a standard measure of inequality that quantifies disparities in exposure by calculating how concentrated exposure is across a set of users. A Gini coefficient close to 1 indicates high inequality (i.e., a few users receive most of the exposure), while a value near 0 suggests a more equal distribution. To measure a user’s exposure within a timeline, we introduce a metric called “weighted occurrence per 1,000 tweets,” defined as the number of times a user’s tweets appear per 1,000 tweets in the timeline, weighted by each tweet’s visibility according to its rank using an exponential decay function—this correction is introduced to give more weight to tweets that appear earlier in one’s timeline, as those tweets are also the more likely to be seen by a user and are known to generate more engagements (Kang and Lerman, 2015). Detailed descriptions of the Gini coefficient calculation and the exposure metric are provided in the Methods section. For each sock-puppet monitoring account, we compute its Gini coefficient with respect to all recommended users in that account’s timelines.
Figure 1. Distribution of Gini coefficient across different groups of accounts. Significant disparities are found in all pairwise comparisons (Mann-Whitney U test: $p$ < 0.001), with right-leaning users experiencing the highest out-of-network exposure inequality.
Figure 1 presents the distribution of Gini coefficients across different account groups: Left-Leaning, Right-Leaning, Balanced, and Neutral. The Mann-Whitney U test reveals that the differences in Gini coefficients between all pairs of groups are significant at the 0.001 level, underscoring meaningful disparities in exposure inequality across these groups. As illustrated in Figure 1, the average Gini coefficient across all groups exceeds 0.45, which suggests a moderate to high level of inequality in exposure on the $\mathbb{X}$ platform. This indicates that algorithmic exposure is somewhat concentrated among certain users rather than evenly distributed.
Notably, right-leaning users experience the highest exposure inequality, followed by balanced and left-leaning users. This suggests that the algorithm’s out-of-network tweet recommendations for right-leaning users are more centralized, reflecting a stronger popularity bias, where a few users dominate exposure. In contrast, neutral users—who do not follow anyone—receive the most diverse recommendations, potentially due to algorithmic cold start, i.e., the absence of information about user preferences that typically informs recommendations (Yuan and Hernandez, 2023).
Since neutral accounts are critical for detecting bias, we took particular care in their setup to ensure neutrality. Neutral accounts follow no other accounts and therefore receive exclusively out-of-network recommendations. This configuration limits any bias that could arise from follow choices, aiming to capture a baseline view of how the algorithm behaves when no user-preferences are specified. However, it is worth noting that certain factors, such as $\mathbb{X}$’s default settings or trending topics, could still introduce slight biases into these recommendations.
Figure 2. Top 20 recommended users in neutral accounts, ranked by their average weighted occurrence per 1,000 tweets. Each box in the boxplot shows the distribution of exposure across all neutral accounts, with red and blue colors indicating right- and left-leaning users, respectively. The figure suggests that right-leaning users are more frequently recommended than left-leaning users in the algorithm’s out-of-network recommendations for neutral accounts.
Now that we understand that out-of-network exposures are skewed toward certain users, an important question arises: Who are these users? Here, we are particularly interested in neutral accounts, which provide an unbiased look at the algorithm’s default behavior. Figure 2 displays the top 20 recommended users in neutral accounts, ranked by their weighted occurrence per 1,000 tweets. Each box in the boxplot represents the distribution of this exposure metric across all neutral accounts. Boxes are colored red or blue to indicate whether the user is right- or left-leaning, based on publicly available data. A qualitative inspection reveals that right-leaning users appear more frequently among the top recommendations than left-leaning users. To quantify this difference, we use the “weighted occurrences per 1,000 tweets” metric: among the top 20 recommended users, right-leaning users make up 30.16% of exposure, compared to 29.34% for left-leaning users. This disparity persists as we expand the pool, with right-leaning users making up 35.26% of exposure in the top 50 (versus 22.34% for left-leaning users) and 31.39% in the top 100 (versus 20.83% for left-leaning users). Further investigations to characterize these users will be conducted in future studies.
In the Appendix, interested readers can find the top 20 recommendations for the left-leaning, right-leaning, and balanced account groups, shedding light on the most amplified accounts across different user groups. A detailed table describing these users’ public information is also provided.
Differential (De-)Amplification of Political Content Among Partisan Accounts
Selective exposure is a psychological concept that refers to the tendency of individuals to prefer information that aligns with their pre-existing beliefs, attitudes, or preferences, while avoiding information that contradicts them (Guess et al., 2018). Algorithms on social media platforms can amplify this effect by recommending content similar to what users already prefer or agree with, reinforcing selective exposure through personalization (Knudsen, 2023). To assess the amplification of certain accounts in partisan users’ timelines, we introduce the “mean amplification ratio” metric, inspired by Huszár et al. (2022). This metric compares the exposure of a user in left-leaning or right-leaning timelines relative to a baseline observed in politically balanced users. A positive mean amplification ratio indicates amplification, while a negative ratio indicates de-amplification. For a detailed description of this metric, please refer to the Methods section.
Figure 3. Amplification ratio of the top 50 recommended users in left-leaning (top) and right-leaning (bottom) accounts, compared to the baseline of balanced accounts. Colored bars indicate a significant difference in exposure metrics (weighted occurrence per 1,000 tweets) between the groups at the 0.05 significance level (using the Mann-Whitney U test), while gray bars indicate no significant difference.
Figure 3 shows the amplification ratio of the top 50 recommended users in left-leaning and right-leaning accounts, compared to a baseline of balanced accounts. Colored bars indicate a significant difference in exposure metrics (weighted occurrence per 1,000 tweets) between groups at the 0.05 significance level (using the Mann-Whitney U test), while gray bars indicate no significant difference. Detailed information about these users is provided in the Appendix. A qualitative inspection reveals that left-leaning sock-puppet accounts tend to see left-leaning users amplified, and right-leaning users de-amplified, with the opposite pattern observed for right-leaning accounts.
For instance, in left-leaning accounts, the top three amplified users are POTUS (the official account of the U.S. President, currently Joe Biden of the Democratic Party), Ron Filipkowski (a former federal prosecutor known for his criticisms of conservative figures), and Mueller, She Wrote (a political commentary and investigative journalism account with a liberal stance). In contrast, the most de-amplified accounts are catturd2 (a right-wing influencer known for political satire), Jack Posobiec (a right-wing media personality and political activist), and Elon Musk (CEO of Twitter/$\mathbb{X}$, who has recently shared conservative viewpoints). This suggests that left-leaning timelines prioritize left-aligned figures while downplaying right-leaning accounts.
On the contrary, for right-leaning accounts, the top three accounts with the highest amplification in right-leaning timelines are atensnut (a conservative commentator), catturd2 (a right-wing influencer known for political satire), and Gunther Eagleman (a conservative content creator). Conversely, the most de-amplified accounts are Rick Wilson (a co-founder of The Lincoln Project), harris_wins (an online community supporting presidential candidate Kamala Harris), and acnewsitics (a liberal-leaning news commentator). This pattern highlights the algorithm’s tendency to amplify conservative figures more heavily in right-leaning timelines while reducing exposure to left-leaning accounts.
Notably, we observe that amplified users in left-leaning group experience a significantly higher magnitude of amplification compared to those in right-leaning group ($M_{left}=41.86\%$, $M_{right}=27.99\%$, Mann-Whitney U $p<0.001$). However, there are no significant differences in the extent of de-amplification between the two groups.
Discussion
In this study, we present preliminary findings from an ongoing audit of algorithmic recommendations on $\mathbb{X}$’s “For You” timelines. Using 120 sock-puppet accounts with left-leaning, right-leaning, balanced, and neutral political orientations, we observe that $\mathbb{X}$ skews exposure toward a select few high-popularity accounts for all users, with right-leaning users experiencing the highest level of inequality. Both left- and right-leaning users see amplified exposure to accounts aligned with their political stance, while exposure to opposing viewpoints is reduced. Additionally, qualitative analysis of neutral accounts with no follow activity reveals a default right-leaning bias in the platform’s recommendations.
Our analysis of exposure inequality aligns with previous studies on algorithmic bias, which have reported similar amplification patterns within users’ in-network content (Bartley et al., 2023). However, our findings diverge from earlier research suggesting that personalized recommendations tend to be more centrist in political stance (Chen et al., 2020; Bouchaud et al., 2023). This discrepancy perhaps highlights a shift in $\mathbb{X}$’s algorithmic behavior, which might have moved away from promoting moderate content to reinforcing users’ existing preferences more explicitly, especially in out-of-network recommendations. The results also add to the growing body of literature indicating that right-leaning accounts are often more prominently featured in algorithmic recommendations, a trend seen here in the default bias toward right-leaning content for new or neutral accounts.
These findings have important implications for the 2024 U.S. Presidential Election. Social media platforms like $\mathbb{X}$ play a significant role in shaping political discourse, and the algorithmic amplification of politically aligned content may influence user perceptions and potentially reinforce echo chambers. The observed right-leaning bias for neutral accounts suggests that users new to the platform or those without strong pre-existing preferences may be more exposed to conservative viewpoints. This could subtly affect user perspectives, especially for those less routinely engaged with political content or discourse.
Another key observation is that, unlike prior research, which has primarily examined the amplification of tweets from media outlets and political figures—especially elected legislators from major political parties (Huszár et al., 2022)—our findings reveal that $\mathbb{X}$’s algorithm now also amplifies political commentators and influencers. This trend is most pronounced in the recommendations for neutral accounts, suggesting a shift in the algorithm’s prioritization toward these types of voices. This shift could be influenced by recent claims that $\mathbb{X}$ prioritizes verified and paid subscription accounts, potentially amplifying influencers who invest in these platform features. The prominence of these non-institutional voices in political content raises questions about the influence of individual commentators on public opinion, as their perspectives may carry a more personal or sensational tone compared to traditional media sources. Adding to the concerns, recent investigations uncovered state-sponsored foreign interference operations with financial backing of prominent political influencers. [^2] Justice Department Disrupts Covert Russian Government-Sponsored Foreign Malign Influence Operation Targeting Audiences in the United States and Elsewhere https://www.justice.gov/opa/pr/justice-department-disrupts-covert-russian-government-sponsored-foreign-malign-influence This underscores the need for further examination into how the recommendation algorithm’s priorities may shape political engagement and public discourse, especially during critical periods like an election year.
In light of these preliminary findings, we emphasize the need for continuous monitoring of algorithmic shifts and their impact on content exposure dynamics. As social media platforms evolve and adjust their algorithms, regular audits are essential to ensure transparency and accountability, particularly when the stakes involve democratic processes and election integrity.
Methods
Data Collection and Dataset Statistics
Data collection for neutral monitoring accounts began around October 2, 2024, and reached a stable deployment of approximately 30 active neutral accounts per day on October 11. Left-leaning, right-leaning, and balanced accounts began appearing consistently in the dataset around October 7, with each group reaching a stable count of about 30 active accounts per day shortly thereafter. Each neutral account receives approximately 500 tweets per session, while each left-leaning, right-leaning, and balanced account receives around 700 tweets per session. Figures 4 display the number of active accounts and the total tweets collected daily.
Exposure Evaluation Metric
For each $\mathbb{X}$ user whose tweet appears in the personalized timelines of our monitoring accounts, we assign a metric called the “weighted occurrence per 1,000 tweets.” This metric is mathematically expressed as:

### v3

# Auditing Political Exposure Bias: Algorithmic Amplification on Twitter/$\mathbb{X}$ During the 2024 U.S. Presidential Election

@@
Abstract.
@@
Approximately 50% of tweets in $\mathbb{X}$’s user timelines are personalized recommendations from accounts they do not follow. This raises a critical question: What political content are users exposed to beyond their established networks, and what implications does this have for democratic discourse online? In this paper, we present a six-week audit of $\mathbb{X}$’s algorithmic content recommendations during the 2024 U.S. Presidential Election by deploying 120 sock-puppet monitoring accounts to capture tweets from their personalized “For You” timelines. Our objective is to quantify out-of-network content exposure for right- and left-leaning user profiles and assess any potential inequalities and biases in political exposure. Our findings indicate that $\mathbb{X}$’s algorithm skews exposure toward a few high-popularity accounts across all users, with right-leaning users experiencing the highest level of exposure inequality. Both left- and right-leaning users encounter amplified exposure to accounts aligned with their own political views and reduced exposure to opposing viewpoints. Additionally, we observe that new accounts experience a right-leaning bias in exposure within their default timelines. Our work contributes to understanding how content recommendation systems may induce and reinforce biases while exacerbating vulnerabilities among politically polarized user groups. We underscore the importance of transparency-aware algorithms in addressing critical issues such as safeguarding election integrity and fostering a more informed digital public sphere.
Algorithmic bias, Social media auditing, Content recommendation systems, Politics, U.S. Presidential Election, Twitter, X
††copyright: acmlicensed††journalyear: 2025††copyright: acmlicensed††conference: Anonymous Conference; Month Day–Day, 2025; Location††booktitle: Proceedings of the Anonymous Conference, Month Day–Day, 2025, Location††ccs: Human-centered computing Empirical studies in collaborative and social computing††ccs: Information systems Social networks
1. Introduction
@@
During the 2024 U.S. Presidential Election, social media platforms like $\mathbb{X}$ (formerly Twitter) play a pivotal role as hubs for political information and public discourse. However, the information users encounter on $\mathbb{X}$ is increasingly curated by algorithmic recommendation systems that personalize content in their “For You” timelines. As of this writing, $\mathbb{X}$’s “For You” timeline typically consists of 50% in-network tweets (i.e., from accounts a given user follows) and 50% out-of-network tweets (i.e., from accounts that user does not directly follow)—an increase from the 37% out-of-network tweets in 2023 (Bouchaud et al., 2023).
How does $\mathbb{X}$’s algorithm select relevant tweets from outside a user’s network? In 2023, Twitter partially open-sourced its recommendation algorithm, revealing that out-of-network recommendations are sourced through engagement and follow graphs, ranked by a neural network, and refined with heuristics and filters (Twitter, Inc., 2023). Despite the increasing prominence of out-of-network tweets in user timelines, much remains unknown about their composition and nature. While prior research has demonstrated amplification of certain political groups and media sources within users’ in-network tweets, the extent to which such biases extend to out-of-network recommendations is unclear. In contexts like the 2024 U.S. Election, examining this issue is essential to understanding how algorithms shape the consumption of online political content and influence users’ perspectives.
@@
Research on $\mathbb{X}$’s algorithmic auditing faces a critical challenge in analyzing out-of-network content: While many studies assess amplification by comparing personalized timelines with reverse-chronological timelines as a baseline—where tweets appear in the order they were posted without algorithmic effects (Bartley et al., 2021, 2023; Huszár et al., 2022; Bouchaud et al., 2023; Wang et al., 2024), out-of-network tweets lack a reverse-chronological baseline as users do not follow the authors of those tweets, making it challenging to quantitatively measure exposure bias. To address this limitation, we utilize a “sock-puppet audit,” a study design that deploys artificial user accounts with controlled features to systematically capture and analyze platform recommendations (Bartley et al., 2021; Chen et al., 2020; Bandy and Diakopoulos, 2021; Hosseinmardi et al., 2024). Specifically, we introduce a self-constructed baseline using accounts that follow a politically balanced social media diet, enabling direct comparisons with other manually-created partisan user accounts. This approach is particularly well-suited to studying out-of-network exposure patterns because it allows us to observe algorithmic behavior without the interference of the variations of user behaviors or connections. Furthermore, previous sock-puppet audits on $\mathbb{X}$ are often constrained by small sample sizes (fewer than 10 accounts) and restricted tweet collection in terms of both quantity and frequency (Bandy and Diakopoulos, 2021; Bartley et al., 2021; Chen et al., 2020), limiting the generalization and robustness of their findings. Our audit aim to implement a more comprehensive data collection strategy, enabling us to systematically observe and analyze algorithmic behaviors on a much larger scale.
Contribution of this work
In this study, we deploy 120 sock-puppet accounts distributed into four groups across the political spectrum—left-leaning, right-leaning, balanced, and neutral—collecting a robust dataset of over 9 million tweets over six weeks from October to November 2024. Within this framework, we systematically evaluate potential exposure biases, such as popularity bias and algorithmic (de-)amplification across account groups.
@@
    We quantify algorithmic exposure to out-of-network content for users with varying political alignments during the 2024 U.S. Election through a sock-puppet audit of $\mathbb{X}$’s personalized timelines.
@@
We find that $\mathbb{X}$ skews exposure toward a few high-popularity accounts for all users, with right-leaning users experiencing the most inequality. Both left- and right-leaning users encounter amplified exposure to accounts aligned with their own political stance and reduced exposure to opposing viewpoints. Additionally, neutral accounts who do not follow anybody (akin to a newly-registered user account) show a default right-leaning bias in content exposure. Our findings reveal how content recommendation systems can influence and amplify biases, potentially increasing vulnerabilities within politically polarized user groups. This work underscores the urgent need for transparent algorithms to safeguard the integrity of online discourse and the sovereignty of elections.
2. Background & Research Questions
2.1. Related Work
The impact of algorithmic content curation on political discourse in social media has been a major focus of research and public debate. Previous studies consistently show that $\mathbb{X}$’s algorithm amplifies political biases and prioritizes high-engagement content, including emotionally charged, toxic, and low-credibility information (Huszár et al., 2022; Bouchaud et al., 2023; Bandy and Diakopoulos, 2021; Chen et al., 2020; Bartley et al., 2021; Corsi, 2024). Researchers have used methods including randomized experiments, sock-puppet audits, crowdsourced audits, and observational data to study $\mathbb{X}$’s algorithmic effects. Some have found that Twitter’s algorithms tend to amplify content from right-leaning media sources and politicians more than their left-leaning counterparts (Huszár et al., 2022; Graham and Andrejevic, 2024). Other studies report increased exposure to ideologically aligned friends (Bouchaud et al., 2023; Bandy and Diakopoulos, 2021), but decreased exposure to external links (Wang et al., 2024; Bandy and Diakopoulos, 2021). Studies also observe increased low-credibility content in algorithmic timelines (Corsi, 2024), with right-leaning users experiencing higher exposure to such content (Chen et al., 2020). Although algorithms are often flagged for promoting ideological bias and political polarization (Barberá, 2020), as observed on platforms like YouTube (Haroon et al., 2023), other analyses of $\mathbb{X}$ and YouTube suggest that its algorithm tends to push centrist content to partisan users (Chen et al., 2020; Hosseinmardi et al., 2024) and displays a more diverse political mix overall (Bouchaud et al., 2023; Wang et al., 2024).
Despite these insights, the existing literature has a key limitation. Algorithmic timelines consist of two distinct components: the reordering and filtering of in-network tweets and the rendering of out-of-network recommendations. While most studies treat the timeline as a unified entity, making it difficult to disentangle biases between these components, our study focuses explicitly on the latter—out-of-network recommendations—which has received little attention in prior research. Our focus is particularly relevant after Elon Musk’s takeover of the platform, as subsequent changes to content moderation and algorithmic priorities (Twitter, Inc., 2023) may have heightened the impact of out-of-network recommendations on user experiences. In what follows, we outline the algorithmic biases under investigation and introduce our research questions (RQs).
2.2. Exposure Inequality
@@
One significant aspect of algorithmic biases on social media is popularity bias (Nikolov et al., 2019). Algorithms often tend to amplify content from certain users over others, creating inequalities in exposure (Bartley et al., 2023). For instance, Twitter’s ranking algorithm employs a ~48M parameter neural network, which uses thousands of features to score each tweet based on engagement probabilities, prioritizing content with higher likelihoods of interaction in users’ feeds (Twitter, Inc., 2023). Previous research has shown that popularity biases can lead to a skew in the visibility of tweets when comparing personalized feeds with reverse-chronological ones, and that users are disproportionately exposed to friends’ tweets (Bartley et al., 2023, 2021). Yet, it remains unclear whether exposure inequalities extend beyond friends to include a broader set of recommended users. Specifically, we pose the following RQ:
RQ1: To what extent do personalized recommendations in $\mathbb{X}$ exhibit exposure inequality among users, and how do these inequalities differ based on political leanings?
2.3. Out-of-Network (De-)Amplification
Another key dimension of bias is ideological bias, particularly its relationship with algorithmic (de-)amplification and selective exposure to political content. Selective exposure is a psychological concept that refers to the tendency of individuals to prefer information that aligns with their pre-existing beliefs, attitudes, or preferences, while avoiding information that contradicts them (Guess et al., 2018). Algorithms on social media platforms can amplify this effect by recommending content similar to what users already prefer or agree with, reinforcing selective exposure through personalization (Knudsen, 2023). Existing research has produced mixed findings on this issue. On one hand, Bakshy et al. (2015) report considerable cross-cutting exposure on Facebook, and Wang et al. (2024) find that Twitter/$\mathbb{X}$ provides higher-quality and less ideologically congruent news curation. On the other hand, Haroon et al. (2023) trained sock puppets to represent five ideological positions ranging from left to right, and found that YouTube’s algorithm consistently promotes ideologically aligned content to partisan users. Given the inconsistency in findings and our focus on algorithmic recommendations, we seek to address the following RQ:
RQ2: Which out-of-network users are (de-)amplified in the timelines of left- and right-leaning accounts compared to balanced accounts?
In the next sections, we outline our experimental setup, data collection process, and methodology designed to address our RQs.
3. Methods
3.1. Experimental Setup
We create 120 sock-puppet accounts on $\mathbb{X}$ divided into four groups based on their political leaning: 30 neutral accounts (default setting, following no one), 30 left-leaning accounts, 30 right-leaning accounts, and 30 balanced accounts. One general concern about sock-puppet audits is ecological validity—whether artificial accounts accurately represent real user behavior and interactions (Wang et al., 2024; Bandy and Diakopoulos, 2021; Bandy, 2021). However, since we are auditing political biases in algorithmic recommendations, which can be influenced by user engagement and community affiliation, it is crucial to control user behavior as much as possible. Previous studies often deploy bots that mimic real-world content consumption by replicating real user follows as “preset” (Bandy and Diakopoulos, 2021; Hosseinmardi et al., 2024). However, real users often follow diverse, non-political accounts, which could confound our focus on political contents. To address this, we limit our sock-puppet accounts to follow exclusively media, political figures, and entities.
We define the orientation of these sock-puppets based on the accounts they follow. To categorize the political alignment of accounts to follow, we use the AllSides Media Bias Chart, [^1] AllSides Media Bias Chart https://www.allsides.com/media-bias/media-bias-chart which rates news sources on a spectrum from left to right based on their political bias. Each left-leaning and right-leaning account follows 10 media outlets, including seven outlets with a moderate (center-left or center-right) bias and three with a stronger (left or right) bias, as defined by the AllSides’ chart. This selection ensures that these accounts represent a realistic mix of moderately and strongly aligned sources, enhancing the accuracy of our analysis of political exposure. Additionally, left-leaning accounts follow key Democratic figures and entities (Kamala Harris, Tim Walz, House Democrats, and Senate Democrats), while right-leaning accounts follow their Republican counterparts (Donald Trump, JD Vance, House Republicans, and Senate Republicans). Balanced accounts, designed to reflect a centrist perspective, follow five center-left and center-right media outlets and both presidential candidates from each major party. All media follows are randomly selected from the respective groups in the media bias chart, ensuring consistency with each group’s intended alignment.
While some studies, particularly on YouTube, allow bots to interact with algorithms (e.g., following recommendations to study radicalization (Hosseinmardi et al., 2024; Haroon et al., 2023)), we refrain from inducing interactions in the current study for several reasons. First, interactions can create feedback loops that distort the algorithm’s outputs, making it difficult to isolate baseline biases. Second, interaction-based designs complicate comparisons across accounts, as partisan accounts might engage differently with recommendations, introducing variability that is hard to standardize. Third, our focus is on measuring how algorithms recommend political content based on baseline configurations, such as predefined follows. Unlike radicalization studies, which examine user-algorithm feedback, our goal is to capture inherent biases in the recommendation system, best analyzed without user interactions.
We also take efforts to mitigate bias in the design of sock-puppet accounts. According to the $\mathbb{X}$ platform, each account was required to select at least three interests at the time of creation. We use a program to select these interest randomly, alongside random birthdates between 1990 and 1999. To further randomize account attributes and mitigate location-based biases in recommendations, a VPN was used during data collection. These steps ensured consistent and relatively unbiased data capture while adhering to platform constraints.
3.2. Data Collection
We develop a timeline crawler to systematically collect tweets recommended to different types of user profiles in $\mathbb{X}$’s “For You” timeline. The timelines for each account are collected four times daily, yielding approximately 500–700 tweets per session, or about 2,000–3,000 tweets per account per day, within the limits that $\mathbb{X}$’s terms of service impose on new, non-premium accounts. The choice of four daily scraping sessions was made to capture the variability in recommendations throughout the day, as the content recommended by $\mathbb{X}$’s algorithm can shift based on temporal factors like recent events or trending topics. It provides a more comprehensive picture of the algorithmic exposure that users might experience. Data collection spanned from October 2, 2024, one month before the election, to November 19, 2024, two weeks after the election, yielding a dataset of 9.79 million tweets. Figure 5 in the Appendix display the number of active accounts and the total tweets collected daily.
@@ 3.2. Data Collection
Table 1 provides an overview of the statistics for the collected tweet dataset across different account types. It shows the average proportion of out-of-network tweets that each account type encounters, with neutral accounts seeing exclusively out-of-network content, while the other accounts have 55%-63% of their timelines composed of out-of-network tweets. Additionally, it details the average proportions of retweets, quoted tweets, and promoted tweets observed by each account type.
@@
Table 1. Statistics of the collected dataset (mean values with standard deviations)
@@
Out-of-network tweet 100% 59.23% (7.45) 55.88% (6.66) 62.27% (5.70)
Retweet 0.15% (0.66) 2.93% (1.07) 2.54% (1.35) 2.41% (1.62)
Quoted tweet 1.37% (2.32) 8.67% (2.33) 12.65% (2.90) 11.98% (2.00)
Promoted tweet 1.36% (1.65) 7.43% (0.60) 7.21% (0.68) 7.84% (1.39)
3.3. Exposure Evaluation Metric
@@
To measure a user’s exposure within a timeline, we introduce a metric called “weighted occurrence per 1,000 tweets,” defined as the number of times a user’s tweets appear per 1,000 tweets in the timeline, weighted by each tweet’s visibility according to its rank. This adjustment gives more weight to tweets that appear earlier in one’s timeline, as those tweets are also the more likely to be seen by a user and are known to generate more engagements (Kang and Lerman, 2015). For each $\mathbb{X}$ user whose tweet appears in the personalized timelines of our monitoring accounts, the “weighted occurrence per 1,000 tweets” metric is mathematically expressed as: